OpenStudy (anonymous):

Find the value of x for which \[\log_{2}(5.2^{x}+1),\log_{4}(2^{1-x}+1),1\] forms an A.P.

OpenStudy (shubhamsrg):

question is log( 5.2^x +1) + log 2 = log (2^1-x +1) where all bases are converted to 2 => 5.2^(x+1) +2 = 2^(1-x) +1 =>10.2^(x) + 1 = 2/ 2^x lets 2^x = y it turns to be a quad eqn. solve it..

OpenStudy (anonymous):

how have you you changed the base of second term?

OpenStudy (shubhamsrg):

log a with base b = log a (base c) / log b (base c) i just used this..

OpenStudy (anonymous):

how?

OpenStudy (shubhamsrg):

RHS = log (2^1-x +1) base 4 = log (2^1-x +1) base 2 / log 4 base 2 clearer r?

OpenStudy (shubhamsrg):

am sorry my presentation might not seem nice,,hmm..